The perimeter of a movie screen is 62 meters. The screen is 11 meters tall. How wide is it?
Answer by EdenWolf(517)

The perimeter formula is 2l+2w=P, so plug in what you know, then solve for w, or the width.

2(11)+2w=62
22+2w=62
2w=40
w=20

It is 20 meters wide.